To predict averge salary for 2005? 1991 $ 275000 and 2001 $1470,000
I also need to know the average salary for 2010? by using 2001 and 1991 amounts

Answer by stanbon(75887)

To predict average salary for 2005?
1991 $ 275000
2001 $1470,000
-------------------
You have two points: (0,275000) : (10,1470,000)
slope = (1470000-275000)/(10-1) = -142,222
Intercept: (0,275000)
---
Equation:
salary = -142,222x + 275000 where x is the number of years after 1991
--------------------------------------------------------
I also need to know the average salary for 2010?
S(9) = -142,222*9 + 275,000 = $147,000.00
---------------------------
Cheers,
Stan H.